Required information [The following information applies to the questions displayed below.] Sweeten Company had no jobs in pro1. What was the companys plantwide predetermined overhead rate? Predetermined overhead rate per MH2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calcula3. What was the total manufacturing cost assigned to Job P? Total manufacturing cost4. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ5. What was the total manufacturing cost assigned to Job Q? Total manufacturing cost6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establi8. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold9. What were the companys predetermined overhead rates in the Molding Department and the Fabrication Department? (Round your10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do n11. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations.) Unit product cost13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final ans14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establ15. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

1. Plant-wide predetermined overhead rate = Estimated total manufacturing overhead / Estimated total machine hours

Plant-wide predetermined overhead rate = $27,800+(2,500*$2.1)+(1,500*$2.9) / 4,000

Plant-wide predetermined overhead rate = $27,800+5,250+4,350 / 4,000 = $9.35 per machine hour

2. Manufcaturing overhead applied::

Job P = 3,700*$9.35 = $34,595

Job Q = 3,100*$9.35 = $28,985

3. Total manufacturing cost assigned to Job P = Direct material+Direct labor+Manufacturing overhead applied

Total manufacturing cost assigned to Job P = $20,000+26,600+34,595 = $81,195

4. Unit product cost Job P = Total manufacturing cost / Units produced

Unit product cost Job P = $81,195 / 20 = $4,060

5. Total manufacturing cost assigned to Job Q = $11,500+10,300+28,985 = $50,785

6. Unit product cost Job Q = $50,785 / 30 = $1,693

7.

Job P Job Q
Total manufacturing cost $81,195 $50,785
Add: Markup@80% 64,956 40,628
Selling price 146,151 91,413
Units produced 20 30
Selling price per unit $7,308 $3,047

8.

Sweeten company cost of goods sold = $81,195+50,785 = $131,980

9. Predetermined overhead = Estimated total manufacturing overhead / Estimated machine hours

Molding department = $11,750+2,500*2.10 / 2,500 = $6.8 per machine hour

Fabrication department = $16,050+1,500*$2.90 / 1,500 = $13.6 per machine hour

10.

Molding department to Job P = 2,400*$6.8 = $16,320

Molding department to Job Q = 1,500*$6.8 = $10,200

11.

Fabrication department to Job P = 1,300*$13.6 = $17,680

Fabrication department to Job Q = 1,600*$13.6 = $21,760

12.

Job P unit product cost = Direct material+Direct labor+Manufacturing overhead applied / Units

Job P unit product cost = $20,000+26,600+16,320+17,680 / 20 = $4,030 per unit

13.

Job Q unit product cost = $11,500+10,300+10,200+21,760 / 30 = $1,792 per unit

14.

Job P Job Q
Total manufacturing cost $80,600 $53,760
Add: Markup@80% 64,480 43,008
Selling price 145,080 96,768
Units produced 20 30
Selling per unit $7,254 $3,226

15.

Sweeten company cost of goods sold = $80,600+53,760 = $134,360